# On discount functions for economic model predictive control without terminal conditions

### v1

@@ 3 Main result
The main result of this work is that a discounted E-MPC scheme without terminal conditions achieves asymptotic stability of any arbitrarily small neighborhood of the optimal orbit $\Pi^{\star}$ if the prediction horizon is chosen sufficiently large and if the discount satisfies Assumption 5.
@@
Let Assumption 1-5 hold. Then, there exists $\varepsilon\in\mathcal{L}$ such that the optimal periodic orbit $\Pi^{\star}$ is practically asymptotically stable w.r.t. $\varepsilon(N)$ under the MPC feedback laws $\mu_{N}^{\beta}$ for all $N\in\mathbb{N}$, i.e., there exists $\gamma\in\mathcal{KL}$ such that for all $x\in\mathbb{X}$, $k\in\mathbb{N}$ it holds that
$\big{\lVert}(x_{\mu_{N}^{\beta}}(k,x),\mu_{N}^{\beta}(x_{\mu_{N}^{\beta}}(k,x)))\big{\rVert}_{\Pi}\leq\max\{\gamma(\lVert x\rVert_{\Pi_{\mathbb{X}}},k),\varepsilon(N)\}.$
@@
We prove the statement by showing that the rotated value function $\tilde{V}_{N}^{\beta}$ is a practical Lyapunov function in the sense of [15, Definition 13]. Showing positive definiteness of $\tilde{\ell}$ and $\tilde{V}_{N}^{\beta}$ is not different from the linearly discounted case [15, Lemma 27, Lemma 28] and is therefore not shown again here. The challenging part is to show that there exists $\delta\in\mathcal{L}$ such that the Lyapunov decrease condition
@@
holds for all $x\in\mathbb{X}$. To establish this inequality, note that $x_{x}^{\star}(1)=x_{\mu_{N}^{\beta}}(1,x)$ and let us apply Lemma A.5 and Lemma A.7 from the Appendix which provide
$\tilde{V}_{N}^{\beta}\left(x_{x}^{\star}(1)\right)-\tilde{V}_{N}^{\beta}(x)$ (18)${\leq}\lambda(x_{x}^{\star}(1))-\lambda(x)+V_{N}^{\beta}(x_{x}^{\star}(1))-V_{N}^{\beta}(x)+\delta_{3}(N)$
(15)${\leq}\lambda(x_{x}^{\star}(1))-\lambda(x)-\ell(x,u_{x}^{\star}(0))+\ell^{\star}+\delta_{2}(N)+\delta_{3}(N)$
=~$=-\tilde{\ell}(x,u_{x}^{\star}(0))+\delta(N)$
with $\delta:=\delta_{2}+\delta_{3}\in\mathcal{L}$. Since $\mu_{N}^{\beta}(x)=u_{x}^{\star}(0)$, this establishes (9) and hence [15, Theorem 14] guarantees practical asymptotic stability with respect to $\varepsilon(N)$ for some $\varepsilon\in\mathcal{L}$. ∎
@@
Proof 3.4.
As $\ell$ is continuous, this is a direct consequence of Theorem 3.1.
@@
In [15], the same asymptotic average performance guarantee was given without assuming minimality of $\Pi^{\star}$ and continuity of $\lambda$. This required an individual proof, as it could not be stated as corollary of the practical asymptotic stability result, which assumed minimality of $\Pi^{\star}$ and continuity of $\lambda$. To keep this paper short, we did not present such a proof but we conjecture that the performance bound could similarly be established without these assumptions.
@@ 4 Numerical Example
Consider the scalar system $x(t+1)=-x(t)+u(t)$ equipped with the stage cost $\ell(x,u)=x^{3}$ and the constraint sets $\mathbb{X}=[-1,1]$ and $\mathbb{U}=[-0.1,0.1]$. The system is optimally operated at the $2$-periodic orbit ${\Pi^{\star}}(0)=(-1,-0.1)$, ${\Pi^{\star}}(1)=(0.9,-0.1)$ with the average cost $\ell^{\star}=\frac{1}{2}(0.9^{3}-1)\approx-0.136$. In Fig. 2, we compare the asymptotic average performance of the proposed discounted E-MPC scheme, starting from the initial condition $x_{0}=0.05$ for the different discount functions $\beta\in\{\beta^{\mathrm{lin}},\beta^{\mathrm{half-lin}},\beta^{\mathrm{non}},\beta^{\mathrm{poly}}\}$ from Example 2.4, where we chose degree $q=2$ for the polynomial discount. We observe that the non-discounted scheme does not converge to the optimal average performance for all prediction horizon lengths $N\neq 3$. For all odd prediction horizons $N\geq 9$ the scheme gets stuck in the steady state $(0.05,0.1)$ and for all even prediction horizons $N\geq 10$ the scheme gets stuck in the periodic orbit $\Pi_{s}(0)=(0.05,-0.1)$, $\Pi_{s}(1)=(-0.15,-0.1)$. This failure is not surprising, as $\beta^{\text{non}}$ does not satisfy Assumption 5 and hence, our theoretical guarantees do not hold for $\beta^{\text{non}}$. Among the other discounts, which satisfy Assumption 5, we can see that $\beta^{\text{half-lin}}$ performs best and achieves optimal performance for all $N\geq 7$, with $\beta^{\text{poly}}$ we need $N\geq 9$, and with $\beta^{\text{lin}}$ we need $N\geq 11$.
@@
For comparison, we take a look at the $p^{\star}$-step E-MPC scheme from [13], which converges to the optimal orbit for all prediction horizon lengths $N\geq 3$ in this example. Hence, in this aspect it is better than the three discounted E-MPC schemes, however, it needs to know the optimal period length $p^{\star}$ in advance, provides only convergence but no stability guarantees, and as discussed in [15, Example 18] may lead to suboptimal transient performance.
Moreover, we compare to an E-MPC with terminal equality constraints (a special case of [14]), i.e., at time $t$ we solve (8) with $\beta=\beta^{\text{non}}$ and the additional constraint that $x_{u}(N,x)={\Pi^{\star}_{\mathbb{X}}}([\phi_{\text{end}}+t]_{2})$ with phase $\phi_{\text{end}}\in\{0,1\}$. We denote the resulting time-dependent MPC feedback by $\mu_{N}^{\mathrm{TC}}(x,t)$. If we start at $t=0$ at $x_{0}={\Pi^{\star}_{\mathbb{X}}}(\phi_{0})$ with phase $\phi_{0}$, then either $[\phi_{0}+N]_{2}=\phi_{\text{end}}$ or $[\phi_{0}+N]_{2}\neq\phi_{\text{end}}$. In the first case, the OCP is feasible for all horizon lengths $N\geq 1$ and the closed-loop follows the optimal orbit ${\Pi^{\star}}$. In the second case, however, we need $N\geq 20$ for feasibility in order to leave $\Pi^{\star}$ and approach it with the synchronized phase again, which additionally leads to a rather high transient cost. For example, starting at $\phi_{0}=0$, i.e., $x_{0}=-1$, choosing $N=20$ and $\phi_{\text{end}}=1$, leads to $J_{20}(x_{0},\mu_{20}^{\mathrm{TC}})=-1$ and $J_{21}(x,\mu_{20}^{\mathrm{TC}})=-0.271$ in closed loop, whereas $J_{20}(x_{0},\mu_{7}^{\beta^{\mathrm{half-lin}}})=-2.71$ and $J_{21}(x,\mu_{7}^{\beta^{\mathrm{half-lin}}})=-3.71$.
@@
Consider the scalar system $x(t+1)=u(t)$, which is equipped with the stage cost $\ell(x,\,u)=-\log(5x^{0.34}-u)$, constrained to the sets $\mathbb{X}=[0,10]$, $\mathbb{U}=[0.1,10]$, and initialized at $x(0)=0.1$. It is known that the optimal cost $\ell^{\star}\approx-1.47$ is achieved at the steady state $\Pi^{\star}\approx(2.23,2.23)$, i.e., we have an optimal orbit of length $p^{\star}=1$ (compare [7, Example 5.1]). Therefore, although $\beta^{\text{non}}$ does not satisfy Assumption 5, we still have guaranteed practical stability from [7]. In Fig. 3, we compare the asymptotic average performance for the discount functions $\beta^{\mathrm{non}}$, $\beta^{\mathrm{lin}}$, $\beta^{\mathrm{half-lin}}$, and $\beta^{\mathrm{poly}}$ with degree $q=2$. As known from [7] and [15, Example 19], we can see that the performance converges exponentially in $N$ for $\beta^{\mathrm{non}}$ and with the approximate rate $\sim\frac{1}{N^{2}}$ for $\beta^{\mathrm{lin}}$. Further, we can see that $\beta^{\mathrm{half-lin}}$ is able to restore the exponential convergence rate and thus, significantly outperforms $\beta^{\mathrm{lin}}$ and $\beta^{\mathrm{poly}}$. For example, for an error of less than $10^{-10}$ we need $N\geq 11$ with $\beta^{\text{non}}$, $N\geq 17$ with $\beta^{\text{half-lin}}$ and $N\geq 6\cdot 10^{4}$ for $\beta^{\text{lin}}$ (compare [15]).
@@ 5 Conclusion and outlook
We have presented in this paper a discounted E-MPC scheme without terminal conditions for nonlinear systems which behave optimally on a periodic orbit. We provided requirements on discount functions such that practical asymptotic stability and asymptotic average performance can be guaranteed. As seen in Example 4.2, the quantitative performance level can vary strongly for different discounts and the performance of the linear discount from [15] can be outperformed significantly. Based on the findings in our numerical examples we conjecture that $\beta^{\mathrm{half-lin}}$ is generally a good choice, i.e., linearly discounting only the second half of the prediction horizon. However, a more extensive numerical study is required to draw conclusions about which discount is best in which situation. Additionally, it is interesting to theoretically investigate under which assumption on the discount function an exponential convergence rate is guaranteed, as is the case for optimal steady-state operation without discounting [7].
Appendix A Technical Lemmas
This appendix states and proves the Lemmas that are needed for the main result.
@@ Appendix A Technical Lemmas
Let Assumption 1-5 hold. We define the number of points of the optimal trajectory $u_{x}^{\star}$ in an $\varepsilon$-neighborhood of the optimal orbit $\Pi^{\star}$ as
$Q_{\varepsilon}^{\beta}(N,\,x):=\#\left\{k\in\mathbb{I}_{[0,N-1]}\big{|}\big{\lVert}\left(x_{x}^{\star}(k),u_{x}^{\star}(k)\right)\big{\rVert}_{\Pi^{\star}}\leq\varepsilon\right\},$ (10)
[…]
Then, $\exists\,\alpha\in\mathcal{K}_{\infty}$ such that
$Q_{\varepsilon}^{\beta}(N,\,x)\geq N-\textstyle\frac{\sqrt{N}}{\alpha(\varepsilon)}$ (11)
holds for all $x\in\mathbb{X}$, $N\in\mathbb{N}$, and $\varepsilon>0$.
@@
First, let us establish an upper bound on the rotated cost functional $\tilde{J}_{N}^{\beta}(x,u_{x}^{\star})=\sum_{k=0}^{N-1}\beta\left(\frac{k}{N}\right)\tilde{\ell}(x_{x}^{\star}(k),u_{x}^{\star}(k))$ starting with
$\tilde{J}_{N}^{\beta}(x,u_{x}^{\star})$ =(6)${=}V_{N}^{\beta}(x)-\sum_{k=0}^{N-1}{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\ell^{\star}+\lambda(x)-\sum_{k=1}^{N}\left({\textstyle\beta\big{(}\frac{k-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\right)\lambda(x_{x}^{\star}(k)).$ ) .
Due to Assumptions 12, and 5 we have
$-\sum_{k=1}^{N}\left({\textstyle\beta\big{(}\frac{k-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\right)\lambda(x_{x}^{\star}(k))$ $\leq\sum_{k=1}^{N}\left({\textstyle\beta\big{(}\frac{k-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\right)\bar{\lambda}=\underbrace{\left(\beta(0)-\beta(1)\right)}_{=1}\bar{\lambda}.$
Due to Assumptions 7 and 4 we can reach the optimal orbit ${\Pi^{\star}_{\mathbb{X}}}$ from $x$ in at most $M_{3}=M_{1}+M_{2}$ steps. Therefore, Lemma 22 from [15] yields
$V_{N}^{\beta}(x)-\sum_{k=0}^{N-1}{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\ell^{\star}+\lambda(x)+\bar{\lambda}\leq M_{3}(\bar{\ell}-\ell^{\star})+2\bar{\lambda}+\ell^{\star}p^{\star}=:C.$
Although this Lemma was only stated for linear discounts, it is straightforward to see that the proof is valid for general $\beta$ satisfying Assumption 5 as only the fact that $\beta$ is non-increasing and $\beta(1)=0$, $\beta(0)=1$ is used. Furthermore, due to Assumption 5 we know that $\mathrm{lim\,inf}_{\xi\to 1}\beta^{\prime}(\xi)>0$ and $\beta$ is non-incresing, hence there must exist a constant $c>0$ such that $\beta$ is lower bounded by the scaled linear discount $\beta(\xi)\geq c\beta^{\text{lin}}(\xi)$ for all $\xi\in[0,1]$. Hence, as $\tilde{\ell}$ is non-negative, we obtain $\tilde{J}_{N}^{\beta^{\text{lin}}}(x,u_{x}^{\star})\leq c^{-1}\tilde{J}_{N}^{\beta}(x,u_{x}^{\star})\leq c^{-1}C$. Therefore, we can apply Lemma 23 from [15], which guarantees that the weak turnpike property holds with $\alpha(\varepsilon)=\sqrt{\frac{c}{2C}\underline{\alpha}_{\tilde{\ell}}(\varepsilon)}$, $\alpha\in\mathcal{K}_{\infty}$. ∎
@@
Let Assumption 15 hold and let $g:\mathbb{X}\times\mathbb{U}\to\mathbb{R}$ be a continuous function and let $g^{\star}:=\frac{1}{p^{\star}}\sum_{k=0}^{p^{\star}-1}g({\Pi^{\star}}(k))$, then there exists $\delta_{1}\in\mathcal{L}$ such that for all $x\in\mathbb{X}$
$\left|g^{\star}-\sum_{k=1}^{N-1}\left({\textstyle\beta\big{(}\frac{k-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\right)g\left(x_{x}^{\star}(k),u_{x}^{\star}(k)\right)\right|\leq\delta_{1}(N).$ (12)
@@
Hence, there are at most $4]{N^{3}}$ points of $(x_{x}^{\star},u_{x}^{\star})$ outside the $\sigma(N)$-neighborhood of ${\Pi^{\star}}$. Denote the left hand side of (12) by $\Sigma_{1}$ and let us split the sum in $\Sigma_{1}$ over $k\in\mathbb{I}_{[1,N-1]}$ into pieces $k\in\mathcal{I}_{j}=\mathbb{I}_{[jp^{\star}+1,(j+1)p^{\star}]}$ of length $p^{\star}$ for $j\in\mathbb{I}_{[0,\lfloor(N-1)/p^{\star}\rfloor-1]}$ and let $\mathcal{I}_{\lfloor(N-1)/p^{\star}\rfloor}=\mathbb{I}_{[\lfloor(N-1)/p^{\star}\rfloor p^{\star},N-1]}$ contain the remainder. Let $\mathcal{J}$ be the set of indices $j$ for which the piece $\mathcal{I}_{j}$ is not fully inside the $\sigma(N)$-neighborhood of ${\Pi^{\star}}$ and of the remainder piece, hence #$\#\mathcal{J}\leq\sqrt[4]{N^{3}}+1$. Now, using Lipschitz continuity $\left|{\textstyle\beta\big{(}\frac{k-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\right|\leq\frac{L_{\beta}}{N}$, we bound all pieces in $\mathcal{J}$ by

### v2

@@ 3 Main result
The main result of this work is that a discounted E-MPC scheme without terminal conditions asymptotically stabilizes a set that contracts to the optimal orbit $\Pi^{\star}$ with increasing prediction horizons if the discount satisfies Assumption 6.
@@
Let Assumption 1-6 hold. Then, there exists $\varepsilon\in\mathcal{L}$ such that $\Pi^{\star}$ is practically asymptotically stable w.r.t. $\varepsilon(N)$ under the discounted E-MPC feedback $\mu_{N}^{\beta}$ for all $N\in\mathbb{N}$, i.e., there exists $\gamma\in\mathcal{KL}$ such that for all $x\in\mathbb{X}$ and $k\in\mathbb{N}$ we have
$\left\lVert\left(x_{\mu_{N}^{\beta}}(k,x),\mu_{N}^{\beta}\big{(}x_{\mu_{N}^{\beta}}(k,x)\big{)}\right)\right\rVert_{\Pi^{\star}}\leq\max\big{\{}\gamma(\lVert x\rVert_{\Pi_{\mathbb{X}}^{\star}},k),\varepsilon(N)\big{\}}.$
@@
Throughout the proof, we assume without loss of generality that $\ell$ is non-negative (cf. [17, Remark 20]). Analogous to $J_{N}^{\beta}$, $V_{N}^{\beta}$, $u_{x}^{\star}$, and $x_{x}^{\star}$, we define $\tilde{J}_{N}^{\beta}$, $\tilde{V}_{N}^{\beta}$, $\tilde{u}_{x}^{\star}$, and $\tilde{x}_{x}^{\star}$ for the case by replacing $\ell$ by the rotated stage cost $\tilde{\ell}$ from (5). We prove the statement by showing that the rotated value function $\tilde{V}_{N}^{\beta}$ is a practical Lyapunov function in the sense of [17, Definition 13]. Showing positive definiteness of $\tilde{\ell}$ and $\tilde{V}_{N}^{\beta}$ is not different from the linearly discounted case [17, Lemma 27, Lemma 28] and is therefore not shown again here. The challenging part is to show that there exists $\delta\in\mathcal{L}$ such that the Lyapunov decrease condition
@@
holds for all $x\in\mathbb{X}$. To establish this inequality, we need the following two Lemmas, which are proven in the Appendix A.
Lemma 3.3 (Almost optimal candidate solution).
$V_{N}^{\beta}\big{(}x_{x}^{\star}(1)\big{)}-V_{N}^{\beta}(x)\leq-\ell\big{(}x,u_{x}^{\star}(0)\big{)}+\ell^{\star}+\delta_{2}(N).$ (9)
Lemma 3.4 (Difference of value functions).
Let Assumptions 1-6 hold. Then, there exists $\delta_{3}\in\mathcal{L}$ such that for all $x,y\in\mathbb{X}$, $N\in\mathbb{N}$ the following inequality holds
These two Lemmas yield
$\tilde{V}_{N}^{\beta}\big{(}x_{x}^{\star}(1)\big{)}-\tilde{V}_{N}^{\beta}(x)$ (10)${\leq}\lambda\big{(}x_{x}^{\star}(1)\big{)}-\lambda(x)+V_{N}^{\beta}\big{(}x_{x}^{\star}(1)\big{)}-V_{N}^{\beta}(x)+\delta_{3}(N)$
(9)${\leq}\lambda\big{(}x_{x}^{\star}(1)\big{)}-\lambda(x)-\ell\big{(}x,u_{x}^{\star}(0)\big{)}+\ell^{\star}+\delta_{2}(N)+\delta_{3}(N)$
=~$=-\tilde{\ell}\big{(}x,u_{x}^{\star}(0)\big{)}+\delta(N)$
with $\delta:=\delta_{2}+\delta_{3}\in\mathcal{L}$. Since $u_{x}^{\star}(0)=\mu_{N}^{\beta}(x)$ and $x_{x}^{\star}(1)=x_{\mu_{N}^{\beta}}(1,x)$, this establishes (8) and hence [17, Theorem 14] guarantees practical asymptotic stability with respect to $\varepsilon(N)$ for some $\varepsilon\in\mathcal{L}$. ∎
@@
Proof 3.6.
For given $N$ and $x$, let $T_{1}\in\mathbb{N}$ be large enough such that $\gamma(\|x\|_{\Pi^{\star}},T_{1})\leq\varepsilon(N)$. For all $T\geq T_{1}$ we split $J_{T}(x,\mu_{N}^{\beta})$ in the interval $\mathbb{I}_{[0,T_{1}-1]}$ and the interval $\mathbb{I}_{[T_{1},T-1]}$ in which $\big{(}x_{x}^{\star}(k),u_{x}^{\star}(k)\big{)}$ is inside the $\varepsilon(N)$ neighborhood of ${\Pi^{\star}}$ due to Theorem 3.1. We use $\bar{\ell}$ to bound the first interval as well as $\varepsilon_{2}\in\mathcal{L}$ and the bound from [17, Lemma 24] to bound the second interval, i.e.,
$\frac{1}{T}J_{T}(x,\mu_{N}^{\beta})\leq\frac{T_{1}}{T}\bar{\ell}+\frac{T-T_{1}+p^{\star}-1}{T}\ell^{\star}+\varepsilon_{2}(N).$
Now, apply $\limsup_{T\to\infty}$ to both sides to obtain the desired result. ∎
@@
In [17], the same asymptotic average performance guarantee was given without assuming minimality of $\Pi^{\star}$ and continuity of $\lambda$. This required an individual proof, as it could not be stated as corollary of the practical asymptotic stability result, which assumed minimality of $\Pi^{\star}$ and continuity of $\lambda$. To keep this paper concise, we did not present such a proof but we conjecture that the performance bound could similarly be established without these assumptions.
@@ 4 Numerical Example
Consider the scalar system $x(t+1)=-x(t)+u(t)$ equipped with the stage cost $\ell(x,u)=x^{3}$ and the constraint sets $\mathbb{X}=[-1,1]$ and $\mathbb{U}=[-0.1,0.1]$. The system is optimally operated at the $2$-periodic orbit ${\Pi^{\star}}(0)=(-1,-0.1)$, ${\Pi^{\star}}(1)=(0.9,-0.1)$ with the average cost $\ell^{\star}=\frac{1}{2}(0.9^{3}-1)\approx-0.136$. In Fig. 1, we compare the asymptotic average performance of the proposed discounted E-MPC scheme, starting from the initial condition $x_{0}=0.05$ for the different discount functions $\beta\in\{\beta^{\mathrm{lin}},\beta^{\mathrm{half-lin}},\beta^{\mathrm{un}},\beta^{\mathrm{poly}}\}$ from Example 2.4, where we chose degree $q=2$ for the polynomial discount. We observe that the undiscounted scheme does not converge to the optimal average performance for all prediction horizon lengths $N\neq 3$. For all odd prediction horizons $N\geq 9$ the scheme gets stuck in the steady state $(0.05,0.1)$ and for all even prediction horizons $N\geq 10$ the scheme gets stuck in the periodic orbit $\Pi_{s}(0)=(0.05,-0.1)$, $\Pi_{s}(1)=(-0.15,-0.1)$. This failure is not surprising, as $\beta^{\text{un}}$ does not satisfy Assumption 6 and hence, our theoretical guarantees do not hold for $\beta^{\text{un}}$. Among the other discounts, which satisfy Assumption 6, we can see that $\beta^{\text{half-lin}}$ performs best and achieves optimal performance for all $N\geq 7$, with $\beta^{\text{poly}}$ we need $N\geq 9$, and with $\beta^{\text{lin}}$ we need $N\geq 11$.
@@
For comparison, we take a look at the $p^{\star}$-step E-MPC scheme from [14], which converges to the optimal orbit for all prediction horizon lengths $N\geq 3$ in this example. Hence, in this aspect it is better than the three discounted E-MPC schemes, however, it needs to know the optimal period length $p^{\star}$ in advance, provides only convergence but no stability guarantees, and as discussed in [17, Example 18] may lead to suboptimal transient performance.
Moreover, we compare to an E-MPC with terminal equality constraints (a special case of [16]), i.e., at time $t$ we solve (7) with $\beta=\beta^{\text{un}}$ and the additional constraint that $x_{u}(N,x)={\Pi^{\star}_{\mathbb{X}}}([\phi_{\text{end}}+t]_{2})$ with phase $\phi_{\text{end}}\in\{0,1\}$. We denote the resulting time-dependent MPC feedback by $\mu_{N}^{\mathrm{TC}}(x,t)$. If we start at $t=0$ at $x_{0}={\Pi^{\star}_{\mathbb{X}}}(\phi_{0})$ with phase $\phi_{0}$, then either $[\phi_{0}+N]_{2}=\phi_{\text{end}}$ or $[\phi_{0}+N]_{2}\neq\phi_{\text{end}}$. In the first case, the OCP is feasible for all horizon lengths $N\geq 1$ and the closed-loop follows the optimal orbit ${\Pi^{\star}}$. In the second case, however, we need $N\geq 20$ for feasibility in order to leave $\Pi^{\star}$ and approach it with the synchronized phase again, which additionally leads to a rather high transient cost. For example, starting at $\phi_{0}=0$, i.e., $x_{0}=-1$, choosing $N=20$ and $\phi_{\text{end}}=1$, leads to $J_{20}\left(x_{0},\mu_{20}^{\mathrm{TC}}\right)=-1$ and $J_{21}\left(x,\mu_{20}^{\mathrm{TC}}\right)=-0.271$ in closed loop, whereas $J_{20}\left(x_{0},\mu_{7}^{\beta^{\mathrm{half-lin}}}\right)=-2.71$ and $J_{21}\left(x,\mu_{7}^{\beta^{\mathrm{half-lin}}}\right)=-3.71$.
@@
Consider the scalar system $x(t+1)=u(t)$, which is equipped with the stage cost $\ell(x,\,u)=-\log(5x^{0.34}-u)$, constrained to the sets $\mathbb{X}=[0,10]$, $\mathbb{U}=[0.1,10]$, and initialized at $x(0)=0.1$. It is known that the optimal cost $\ell^{\star}\approx-1.47$ is achieved at the steady state $\Pi^{\star}\approx(2.23,2.23)$, i.e., we have an optimal orbit of length $p^{\star}=1$ (compare [8, Example 5.1]). Therefore, although $\beta^{\text{un}}$ does not satisfy Assumption 6, we still have guaranteed practical stability from [8]. In Fig. 2, we compare the asymptotic average performance for the discount functions $\beta^{\mathrm{un}}$, $\beta^{\mathrm{lin}}$, $\beta^{\mathrm{half-lin}}$, and $\beta^{\mathrm{poly}}$ with degree $q=2$. As known from [8] and [17, Example 19], we can see that the performance converges exponentially in $N$ for $\beta^{\mathrm{un}}$ and with the approximate rate $\sim\frac{1}{N^{2}}$ for $\beta^{\mathrm{lin}}$. Further, we can see that $\beta^{\mathrm{half-lin}}$ is able to restore the exponential convergence rate and thus, significantly outperforms $\beta^{\mathrm{lin}}$ and $\beta^{\mathrm{poly}}$. For example, for an error of less than $10^{-10}$ we need $N\geq 11$ with $\beta^{\text{un}}$, $N\geq 17$ with $\beta^{\text{half-lin}}$ and $N\geq 6\cdot 10^{4}$ for $\beta^{\text{lin}}$ (compare [17]).
@@ 5 Conclusion and outlook
We have presented in this paper a discounted E-MPC scheme without terminal conditions for nonlinear systems which behave optimally on a periodic orbit. We provided requirements on discount functions such that practical asymptotic stability and asymptotic average performance can be guaranteed. As seen in Example 4.2, the quantitative performance level can vary strongly for different discounts and the performance of the linear discount from [17] can be outperformed significantly. Based on the findings in our numerical examples we conjecture that $\beta^{\mathrm{half-lin}}$ is generally a good choice when dealing with periodic orbits, i.e., linearly discounting only the second half of the prediction horizon. When dealing with steady states (optimal period length $p^{\star}=1$), we found in our examples no advantage of using discounts instead of the standard undiscounted E-MPC scheme. However, a more extensive numerical study is required to draw conclusions about which discount is best in which situation. Additionally, it is interesting to theoretically investigate under which assumption on the discount function an exponential convergence rate is guaranteed, as is the case for optimal steady-state operation without discounting [8].
Appendix A Proof of Lemmas 3.3 and 3.4
In order to prove Lemmas 3.3 and 3.4, we first show in Lemma A.1 that the weak turnpike property (cf. [17, Definition 7]) holds and we use it in Lemma A.3 to compute the average of a function over turnpike trajectories.
@@
Let Assumptions 1-6 hold. We define the number of points of the optimal trajectory $u_{x}^{\star}$ in an $\varepsilon$-neighborhood of the optimal orbit $\Pi^{\star}$ as $Q_{\varepsilon}^{\beta}(N,\,x):=\#\left\{k\in\mathbb{I}_{[0,N-1]}\big{|}\big{\lVert}\big{(}x_{x}^{\star}(k),u_{x}^{\star}(k)\big{)}\big{\rVert}_{\Pi^{\star}}\leq\varepsilon\right\}$. Then, there exists $\alpha\in\mathcal{K}_{\infty}$ such that for all $x\in\mathbb{X}$, $N\in\mathbb{N}$, and $\varepsilon>0$ we have
$Q_{\varepsilon}^{\beta}(N,\,x)\geq N-\textstyle\frac{\sqrt{N}}{\alpha(\varepsilon)}.$ (11)
@@
First, let us establish an upper bound on the rotated cost functional $\tilde{J}_{N}^{\beta}(x,u_{x}^{\star})=\sum_{k=0}^{N-1}\beta\left(\frac{k}{N}\right)\tilde{\ell}\big{(}x_{x}^{\star}(k),u_{x}^{\star}(k)\big{)}$ starting with
$\tilde{J}_{N}^{\beta}(x,u_{x}^{\star})$ =(5)${=}V_{N}^{\beta}(x)-\sum_{k=0}^{N-1}{\textstyle\beta\left(\frac{k}{N}\right)}\ell^{\star}+\lambda(x)-\sum_{k=1}^{N}\left({\textstyle\beta\left(\frac{k-1}{N}\right)}-{\textstyle\beta\left(\frac{k}{N}\right)}\right)\lambda(x_{x}^{\star}(k)).$ ) .
Due to Assumptions 23, and 6 we have
$-\sum_{k=1}^{N}\left({\textstyle\beta\left(\frac{k-1}{N}\right)}-{\textstyle\beta\left(\frac{k}{N}\right)}\right)\lambda(x_{x}^{\star}(k))$ $\leq\sum_{k=1}^{N}\left({\textstyle\beta\left(\frac{k-1}{N}\right)}-{\textstyle\beta\left(\frac{k}{N}\right)}\right)\bar{\lambda}=\underbrace{\left(\beta(0)-\beta(1)\right)}_{=1}\bar{\lambda}.$
Due to Assumptions 4 and 5 we can reach the optimal orbit ${\Pi^{\star}_{\mathbb{X}}}$ from $x$ in at most $M_{3}=M_{1}+M_{2}$ steps. Therefore, Lemma 22 from [17] yields
$V_{N}^{\beta}(x)-\sum_{k=0}^{N-1}{\textstyle\beta\left(\frac{k}{N}\right)}\ell^{\star}+\lambda(x)+\bar{\lambda}\leq M_{3}(\bar{\ell}-\ell^{\star})+2\bar{\lambda}+\ell^{\star}p^{\star}=:C.$
Although this Lemma was only stated for linear discounts, it is straightforward to see that the proof is valid for general $\beta$ satisfying Assumption 6 as only the fact that $\beta$ is non-increasing and $\beta(1)=0$, $\beta(0)=1$ is used. Furthermore, due to Assumption 6 we know that $\mathrm{lim\,sup}_{\xi\to 1}\beta^{\prime}(\xi)<0$ and $\beta$ is non-incresing, hence there must exist a constant $c>0$ such that $\beta$ is lower bounded by the scaled linear discount $\beta(\xi)\geq c\beta^{\text{lin}}(\xi)$ for all $\xi\in[0,1]$. Hence, as $\tilde{\ell}$ is non-negative, we obtain $\tilde{J}_{N}^{\beta^{\text{lin}}}(x,u_{x}^{\star})\leq c^{-1}\tilde{J}_{N}^{\beta}(x,u_{x}^{\star})\leq c^{-1}C$. Therefore, we can apply Lemma 23 from [17], which guarantees that the weak turnpike property holds with $\alpha(\varepsilon)=\sqrt{\frac{c}{2C}\underline{\alpha}_{\tilde{\ell}}(\varepsilon)}$, $\alpha\in\mathcal{K}_{\infty}$. ∎
@@
Let Assumptions 16 hold and let $g:\mathbb{X}\times\mathbb{U}\to\mathbb{R}$ be a continuous function and let $g^{\star}:=\frac{1}{p^{\star}}\sum_{k=0}^{p^{\star}-1}g({\Pi^{\star}}(k))$, then there exists $\delta_{1}\in\mathcal{L}$ such that for all $x\in\mathbb{X}$
$\left|g^{\star}-\sum_{k=1}^{N-1}\left({\textstyle\beta\left(\frac{k-1}{N}\right)}-{\textstyle\beta\left(\frac{k}{N}\right)}\right)g\left(x_{x}^{\star}(k),u_{x}^{\star}(k)\right)\right|\leq\delta_{1}(N).$ (12)
@@
Hence, there are at most $4]{N^{3}}$ points of $(x_{x}^{\star},u_{x}^{\star})$ outside the $\sigma(N)$-neighborhood of ${\Pi^{\star}}$. Denote the left hand side of (12) by $\Sigma_{1}$ and let us split the sum in $\Sigma_{1}$ over $k\in\mathbb{I}_{[1,N-1]}$ into pieces $k\in\mathcal{I}_{j}=\mathbb{I}_{[jp^{\star}+1,(j+1)p^{\star}]}$ of length $p^{\star}$ for $j\in\mathbb{I}_{[0,\lfloor(N-1)/p^{\star}\rfloor-1]}$ and let $\mathcal{I}_{\lfloor(N-1)/p^{\star}\rfloor}=\mathbb{I}_{[\lfloor(N-1)/p^{\star}\rfloor p^{\star},N-1]}$ contain the remainder. Let $\mathcal{J}$ be the set of indices $j$ for which the piece $\mathcal{I}_{j}$ is not fully inside the $\sigma(N)$-neighborhood of ${\Pi^{\star}}$ and of the remainder piece, hence #$\#\mathcal{J}\leq\sqrt[4]{N^{3}}+1$. Now, using Lipschitz continuity $\left|{\textstyle\beta\left(\frac{k-1}{N}\right)}-{\textstyle\beta\left(\frac{k}{N}\right)}\right|\leq\frac{L_{\beta}}{N}$, we bound all pieces in $\mathcal{J}$ by